Preform the indicated operation:(z-5)(z+2)
Thanks,
Pat:)

Answer by stanbon(48526)

Perform the indicated operation:(z-5)(z+2)
Use FOIL to get:
F: z*z = z^2
O: 2*z = 2z
I: -5*z = -5z
L: -5*2 = -10
-----------
Combine these products to get:
z^2 -3z -10
===============
Cheers,
Stan H.